OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24

OpenStudy (anonymous):

my technique is kind of different from most of the people's.

OpenStudy (anonymous):

I'll write down the steps.

OpenStudy (anonymous):

To begin with, I chose to get rid of x first. so I would multiply by -2 on the first equation and 3 on the second equation. and that becomes -6x+12y=-24 -6x+9y =18

OpenStudy (anonymous):

and you change the signs on the second equation, and that will give you, -6x+12y=-24 6x-9y =-18

OpenStudy (anonymous):

you add them. 3y=-42 y=-14

OpenStudy (anonymous):

now since you got the value of y, you plug in to whatever equation you want to plug in it to. let's plug in the y value to the second equation. (easier numbers) -2x+3(-14)=6 -2x-42=6 -2x=48 x=-24

OpenStudy (anonymous):

oh yeah, they wanted you to find y value that's great! then it's going to y=-14

OpenStudy (anonymous):

Thanks so much! :)

OpenStudy (anonymous):

sure